OpenStudy (anonymous):

The domain of the function f(x)=8x / 5x−1

OpenStudy (anonymous):

pls put in interval notation

OpenStudy (anonymous):

u cant have denominator = 0, so 5x-1 not= 0. that gives us x not= 1/5 domain = (-infinity, 1/5)u(1/5,infinity) or domain = (-infinity,infinity)-{1/5}

OpenStudy (anonymous):

hey thanks for the help. can you help me with this also The domain of the function f(x)=−15x+12 / x^2−7x+10

OpenStudy (anonymous):

factor out the denominator u get (x-5)(x-2) denominator not= 0 so x not=5 and x not=2 if u put into the notation form x = (-infinity,infinity)-{2,5} or x= (-infinty,2) u (2,5) u (5,infinity)

OpenStudy (anonymous):

thanks for the help do you mind, i have more questions?